OpenStudy (anonymous):

For f(x) = 2x + 8, find x when f(x) = -7.

OpenStudy (anonymous):

Plug in -7 or x F(x) = 2x + 8 F(-7) = 2(-7) + 8 F(-7) = -14 + 8 F(-7) = -6 If you want to do one more, I can help you. Tell me if you need help or are having a hard time understanding :)